OpenStudy (anonymous):

Solve the system using substitution. 3x – y = -1 x – y = -3 What is the solution for y? -4 -1 1 4

OpenStudy (anonymous):

3x - y = -1 x - y = -3 x + 3 = y 3x - (x+3) = -1 2x - 3 = -1 2x = 2 x = 1 y = 4

OpenStudy (anonymous):

yeah 4 is the answer